16.7 pts Question 4 Find the value of x in the following rectangles 20 10. 11. 12. 16 15 10) 11) 12)
Answer #1

Soln (10 D a 15 с O ABC According to Pythagorues theorem ( hypotenuse) = (oppositeje+ (Adjacent)? Ac² AB²+ Bc² x² = 198+ (15)Hence n = 17.4928 A 224 u B 20 Since с AD un BC I because ABCD is a rectangle) BL = 20 BCD According to pythagorous theorem B